### v3

@@ Abstract
Mathematical models of protein-protein dynamics, such as the heterodimer model, play a crucial role in understanding many physical phenomena. This model is a system of two semilinear parabolic partial differential equations describing the evolution and mutual interaction of biological species. An example is the neurodegenerative disease progression in some significant pathologies, such as Alzheimer’s and Parkinson’s diseases, characterized by the accumulation and propagation of toxic prionic proteins. This article presents and analyzes a flexible high-order discretization method for the numerical approximation of the heterodimer model. We propose a space discretization based on a Discontinuous Galerkin method on polygonal/polyhedral grids, which provides flexibility in handling complex geometries. Concerning the semi-discrete formulation, we prove stability and a-priori error estimates for the first time. Next, we adopt a $\theta$-method scheme as a time integration scheme. Convergence tests are carried out to demonstrate the theoretical bounds and the ability of the method to approximate traveling wave solutions, considering also complex geometries such as brain sections reconstructed from medical images. Finally, the proposed scheme is tested in a practical test case stemming from neuroscience applications, namely the simulation of the spread of $\alpha$-synuclein in a realistic test case of Parkinson’s disease in a two-dimensional sagittal brain section geometry reconstructed from medical images.
@@ 1 Introduction
The heterodimer continuum model [1] is a well-recognized mathematical model in the context of the description of protein-protein interactions [2, 3]. From a mathematical point of view, this model is a system of parabolic semilinear partial differential equations (PDEs) [4]. Its formulation allows the description of the underlying microscopic mechanism, typical of protein dynamics while capturing the broader macroscopic phenomena of propagation in spatial domains. An important application of this equation is the study of the prionic protein dynamics in neurodegenerative diseases.
A typical model used in literature is the Fisher-Kolmogorov (FK) equation. This is useful in describing neurodegenerative diseases [5], and more in general for biological applications. The latter can be obtained as of the heterodimer model, under the assumption of almost constant dynamics of one of two populations (i.e. healthy protein concentration [6]). However, the more expensive computational effort of the heterodimer model can be justified considering the possibility of analyzing a more informative solution, as it explicitly accommodates the dynamic evolution of both biological populations. However, the mathematical complexity of the formulation is higher, considering two coupled equations whose traveling-wave solution is not guaranteed for all the values of the parameters [1].
@@
In neurodegeneration, considering the pathologies known as proteinopathies (i.e. Alzheimer’s and Parkinson’s diseases), the heterodimer model embraces a complete consideration of both healthy and misfolded protein configurations, typically related to the pathology development [7]. Furthermore, the model includes explicit terms representing distinct molecular kinetic processes: production, clearance [8], and conversion, which is typically guided by the presence of other misfolded proteins [9]. This level of detail of the mathematical model affords the ability to simulate different scenarios [10]. Finally, because these phenomena typically span over decades, clinical observations are very complex and the development of mathematical and computational models can be a key instrument to asses patient-specific evolution scenarios.
@@
Concerning the numerical approximation of the heterodimer model, in literature we find many different methods, i.e. finite element methods [10] or reduced order network diffusion models [11]. For what concerns numerical methods for semilinear parabolic problems there exist some works concerning discontinuous Galerkin approximation for the vectorial case [12, 13, 14], as well as for the scalar case [15, 5, 16, 17]. In this work, we propose and analyze a Discontinuous Galerkin formulation on polygonal/polyhedral grids (PolyDG) for the space discretization of the heterodimer system, coupled with a Crank-Nicolson time discretization scheme, with a semi-implicit treatment of the nonlinear term. For this method, we first derive a stability estimate. The derived estimate is not global in time, due to the treatment of the quadratic nonlinear coupling terms that cannot be simplified by simple choices of the test functions and that need an interpolation using the Gagliardo-Nirenberg-type inequality. Differently from the scalar case of the FK equation in [5], we improve the stability-bound properties by eliminating the exponential dependence on time. Moreover, we prove an a-priori error estimate for the semi-discrete formulation. Also in this case, some difficulties arise compared to the FK model, but by using different interpolation inequalities, we can find an error estimate without a hypothesis on the model parameters, differently from [5]. To the best of our knowledge, these are the first results of this type for the semi-discrete form of the heterodimer model.
@@
The proposed approach enjoys distinguishing features compared to the available approaches in the literature. First, it supports high-order approximations able to correctly reproduce traveling-wave solutions. Leveraging the PolyDG method, we attain flexibility in addressing the complexities of geometry. This is connected to the possibility of using arbitrarily shaped elements, exploiting also the possibility of taking advantage of agglomeration techniques [18]. The procedure consists of constructing a coarse mesh starting from an original highly refined triangulation, needed to accurately represent complex domains, as in the application under consideration. The agglomerated grid now comprises generic polygons, maintaining the original boundary details without requiring curved elements, leading to a noteworthy reduction in computational cost. Finally, PolyDG methods allow local tuning of the discretization parameters, such as polynomial order [19]. This flexibility becomes particularly advantageous in the accurate approximation of propagating wavefronts, even within meshes that are not excessively refined.
@@ 2 The mathematical model
  • $k_{0},k_{1},\tilde{k}_{1},k_{12}\in L^{\infty}_{+}(\Omega)$ and $\exists K_{1}>0:\,K_{1}\leq k_{1}$ and $\exists\tilde{K}_{1}>0:\,\tilde{K}_{1}>\leq\tilde{k}_{1}$
@@
By analyzing the steady-state solutions of (1) with constant forcing term in the first equation $f_{c}(\boldsymbol{x},t)=k_{0}(\boldsymbol{x})$ and null in the second $f_{q}(\boldsymbol{x},t)=0$, it can be proved that the model exhibits two distinct equilibria. The first is an unstable equilibrium at $c=k_{0}/k_{1}$ and $q=0$, representing the extinction of the second population. The second equilibrium $c=\tilde{k}_{1}/k_{12}$ and $q=k_{0}/\tilde{k}_{1}-k_{1}/k_{12}$ is stable and corresponds to a situation where both species coexist. This implies that even a small non-zero quantity of the second population, can initiate an irreversible chain reaction of conversion and spread [24]. Furthermore, for the observation of a travelling-wave solution, transitioning the system towards a coexistence state, it is crucial for the coefficients of Equations (1) to adhere to a specific relationship [11]:
@@ 3.1 Discrete setting: polytopal meshes and functional spaces
We proceed by defining the functional spaces relevant to our analysis. We define the Broken Sobolev space $W_{q}=H^{q}(\Omega,\mathscr{T}_{h})=\{v\in L^{2}(\Omega):\quad v|_{K}\in H^{q}(K)\;\forall K\in\mathscr{T}_{h}\}$. We will seek discrete solutions to the problem within the following discontinuous finite element space:
@@ 3.2 PolyDG semi-discrete formulation
$\gamma_{F}=\gamma_{0}\begin{dcases}\max\left\{\{d^{K}\}_{\mathrm{H}},\{k^{K}\}_{\mathrm{H}}\right\}\dfrac{p^{2}}{\{h\}_{\mathrm{H}}},&\mathrm{on}\;F\in\mathscr{F}^{\mathrm{I}}_{h},\\ \max\left\{d^{K},k^{K}\right\}\dfrac{p^{2}}{h},&\mathrm{on}\;F\in\mathscr{F}_{h}^{\mathrm{D}},\end{dcases}$ (4)
where $\gamma_{0}$ is a constant parameter that should be chosen sufficiently large to ensure the stability of the discrete formulation, $d^{K}=\|\sqrt{\mathbf{D}|_{K}}\|^{2}$, and $k^{K}=\|\,(1+k_{12}|_{K})(k_{1}|_{K}+\tilde{k}_{1}|_{K})\|$. We define for any $u_{h},w_{h},v_{h}\in W_{h}^{\mathrm{DG}}$ the following bilinear and trilinear forms:
@@ 4 Analysis of the semi-discrete PolyDG formulation
$\|c\|_{\mathrm{DG}}=$ $\left\|\sqrt{\mathbf{D}}\nabla_{h}c\right\|+\left\|\sqrt{\gamma_{F}}[\![c]\!]\right\|_{\mathscr{F}^{\mathrm{I}}_{h}\cup\mathscr{F}_{h}^{\mathrm{D}}}\quad\forall c\in W_{1}$ (7a)
$|\!\!\>|\!\!\>|{c}|\!\!\>|\!\!\>|_{\mathrm{DG}}=$ $\|c\|_{\mathrm{DG}}+\left\|\gamma_{F}^{-1/2}\{\!\!\{\mathbf{D}\nabla_{h}c\}\!\!\}\right\|_{\mathscr{F}^{\mathrm{I}}_{h}\cup\mathscr{F}_{h}^{\mathrm{D}}}\forall c\in W_{2}.$ (7b)
We recall that the previous norms are equivalent in the space of discontinuous functions $W_{h}^{\mathrm{DG}}$, under Assumption 2 thanks to the following trace-inverse inequality [29]:
$\|c\|_{L^{2}(\partial K)}\lesssim\frac{p^{2}}{h_{K}}\|c\|_{L^{2}(K)}\quad\forall c\in\mathbb{P}_{p}(K)\quad\forall K\in\mathscr{T}_{h}.$ (8)
@@
$\|c(t)\|_{\varepsilon}^{2}=\|c(t)\|^{2}+\int_{0}^{t}\left(\|c(s)\|_{\mathrm{DG}}^{2}+\|c(s)\|^{2}\right)\mathrm{d}s,\qquad|\!\!\>|\!\!\>|c(t)|\!\!\>|\!\!\>|_{\varepsilon}^{2}=\|c(t)\|+\int_{0}^{t}\left(|\!\!\>|\!\!\>|c(s)|\!\!\>|\!\!\>|_{\mathrm{DG}}^{2}+\|c(s)\|^{2}\right)\mathrm{d}s.$ (9)
@@ 4.1 Preliminary technical results
Let $a,c$ be three positive constants and let $u\in L^{\infty}_{+}(0,\hat{t})=\{u\in L^{\infty}(0,\hat{t}):\;u(t)\geq 0$ a.e. in $(0,\hat{t})\}$ such that:
$u(t)\leq a+c\int_{0}^{t}u^{\eta}(s)\mathrm{d}s,\qquad\mathrm{for\;almost\;any}\;t\in(0,\hat{t}),$ (10)
@@
$u(t)\leq\dfrac{a}{\left(1-a^{\eta-1}\,c\,(\eta-1)t\right)^{\frac{1}{\eta-1}}}.$ (12)
@@ 7 Simulation of $\alpha$-synuclein spreading in a two dimensional brain section
$\alpha=k_{12}\frac{k_{0}}{k_{1}}-\tilde{k}_{1}.$

### v4

@@ Abstract
Mathematical models of protein-protein dynamics, such as the heterodimer model, play a crucial role in understanding many physical phenomena, e.g., the progression of some neurodegenerative diseases. This model is a system of two semilinear parabolic partial differential equations describing the evolution and mutual interaction of biological species. This article presents and analyzes a high-order discretization method for the numerical approximation of the heterodimer model capable of handling complex geometries. In particular, the proposed numerical scheme couples a Discontinuous Galerkin method on polygonal/polyhedral grids for space discretization, with a $\theta$-method for time integration. This work presents novelties and progress with respect to the mathematical literature, as stability and a-priori error analysis for the heterodimer model are carried out for the first time. Several numerical tests are performed, which demonstrate the theoretical convergence rates, and show good performances of the method in approximating traveling wave solutions as well as its flexibility in handling complex geometries. Finally, the proposed scheme is tested in a practical test case stemming from neuroscience applications, namely the simulation of the spread of $\alpha$-synuclein in a realistic test case of Parkinson’s disease in a two-dimensional sagittal brain section geometry reconstructed from medical images.
@@ 1 Introduction
The heterodimer continuum model [1] is a well-recognized mathematical model in the context of the description of protein-protein interactions [2, 3]. From a mathematical point of view, this model is a system of parabolic semilinear partial differential equations (PDEs) [4]. Its formulation allows the description of the underlying microscopic mechanism, typical of protein dynamics, while simultaneously capturing the broader macroscopic phenomena of propagation in spatial domains.
@@
An important application of this model concerns the study of prionic protein dynamics in neurodegenerative diseases. Indeed, in those pathologies known as proteinopathies (i.e. Alzheimer’s and Parkinson’s diseases), the heterodimer model embraces a complete consideration of both healthy and misfolded protein configurations, typically related to the pathology development [5]. Furthermore, the model includes explicit terms representing distinct molecular kinetic processes: production, clearance [6], and conversion, which is typically guided by the presence of other misfolded proteins [7]. This level of detail of the mathematical model affords the ability to simulate different scenarios [8], such as variations in clearance or production mechanisms. The development of mathematical models and computational methods in this field, where the neurodegeneration may span over decades and the clinical data are not easy to collect, represents an essential tool to support medical science and a key instrument to asses patient-specific evolution scenarios. In biological applications, and in particular, in the field of neurodegenerative diseases, a well-known model is the Fisher-Kolmogorov (FK) equation (see [9, 3, 10]). The latter can be interpreted as a particular case of the heterodimer model, when the dynamic of one population is constant (i.e. healthy protein concentration [9]). Instead, the heterodimer model accommodates the dynamics of both biological populations, hence delivering a more informative solution. The price to pay is a substantially increased complexity of the mathematical formulation, which involves not just one equation, but rather a system of two equations, together with a larger computational effort required to approximate the solution of the model.
@@
Concerning the numerical approximation of the heterodimer model, in literature we find many different methods, i.e. finite element methods [8] or reduced order network diffusion models [11]. For what concerns numerical methods for semilinear parabolic problems, there exist some works concerning Discontinuous Galerkin approximation for the vectorial case [12, 13, 14], as well as for the scalar case [15, 10, 16, 17]. In this work, we propose and analyze a Discontinuous Galerkin formulation on polygonal/polyhedral grids (PolyDG) for the space discretization of the heterodimer system, coupled with a Crank-Nicolson time discretization scheme, with a semi-implicit treatment of the nonlinear terms.
@@
The proposed approach enjoys distinguishing features and relevant advantages, when compared to the available approaches in the literature. First, it supports high-order approximations that are able to correctly reproduce traveling-wave solutions. Second, it is capable to handle complex geometries. This is due to the possibility of using arbitrarily shaped polygonal/polyhedral elements, possibly obtained by means of agglomeration techniques [18]. Finally, PolyDG methods allow the local tuning of discretization parameters, like the polynomial order [19]. This flexibility becomes particularly advantageous in the accurate approximation of propagating wavefronts, even within meshes that are not excessively refined.
The main contributions of the present work are the following. On the one hand, we study the PolyDG semi-discrete formulation. The performed analysis contains an important improvement with respect the previous results for the FK equation in [10]: indeed, the exponential dependence on time of the stability bound has been eliminated. Moreover, the analysis is more challenging as it involves quadratic nonlinear coupling terms that must be treated using the Gagliardo-Nirenberg inequality. On the other hand, we prove an a-priori error estimate for the PolyDG semi-discrete formulation. Also in this case, some improvements have been achieved when compared to [10], namely, error estimates are devied without a assumptions on the model parameters. To the best of our knowledge, these are the first results of this type for the semi-discrete form of the heterodimer model.
@@ 2 The mathematical model
  • $k_{1},\tilde{k}_{1},k_{12}\in L^{\infty}_{+}(\Omega)$ and $\exists K_{1}>0:\,K_{1}\leq k_{1}$ and $\exists\tilde{K}_{1}>0:\,\tilde{K}_{1}\leq\tilde{k}_{1}$
@@
By analyzing the steady-state solutions of (1) with constant forcing term in the first equation $f_{c}(\boldsymbol{x},t)=k_{0}$ and null in the second $f_{q}(\boldsymbol{x},t)=0$, it can be proved that the model exhibits two distinct equilibria. The first is an unstable equilibrium at $c=k_{0}/k_{1}$ and $q=0$, representing the extinction of the second population. The second equilibrium $c=\tilde{k}_{1}/k_{12}$ and $q=k_{0}/\tilde{k}_{1}-k_{1}/k_{12}$ is stable and corresponds to a situation where both species coexist. This implies that even a small non-zero quantity of the second population, can initiate an irreversible chain reaction of conversion and spread [24]. Furthermore, for the observation of a traveling-wave solution, transitioning the system towards a coexistence state, it is crucial for the coefficients of Equations (1) to adhere to a specific relationship [11]:
@@ 3.1 Discrete setting: polytopal meshes and functional spaces
We proceed by defining the functional spaces relevant to our analysis. We define the Broken Sobolev space:
$W_{q}=H^{q}(\Omega,\mathscr{T}_{h})=\{v\in L^{2}(\Omega):\quad v|_{K}\in H^{q}(K)\;\forall K\in\mathscr{T}_{h}\}.$ (4)
We will seek discrete solutions to the problem within the following discontinuous finite element space:
@@ 3.2 PolyDG semi-discrete formulation
$\gamma_{F}=\gamma_{0}\begin{dcases}\max\left\{\{d^{K}\}_{\mathrm{H}},\{k^{K}\}_{\mathrm{H}}\right\}\dfrac{p^{2}}{\{h_{K}\}_{\mathrm{H}}},&\mathrm{on}\;F\in\mathscr{F}^{\mathrm{I}}_{h},\\ \max\left\{d^{K},k^{K}\right\}\dfrac{p^{2}}{h_{K}},&\mathrm{on}\;F\in\mathscr{F}_{h}^{\mathrm{D}},\end{dcases}$ (5)
where $\gamma_{0}$ is a constant parameter that should be chosen sufficiently large to ensure the stability of the discrete formulation, $d^{K}=\|\sqrt{\mathbf{D}|_{K}}\|^{2}$, and $k^{K}=\|\,(1+k_{12}|_{K})(k_{1}|_{K}+\tilde{k}_{1}|_{K})\|$. In Equation (5), the symbol {}$\{\cdot\}_{\mathrm{H}}$ denotes the the harmonic average operator, which is defined as $\{\omega^{K}\}_{\mathrm{H}}=\frac{2\omega^{K^{+}}\omega^{K^{-}}}{\omega^{K^{+}}+\omega^{K^{-}}}$ for any function $\omega^{K}$ associated with the partition $\mathscr{T}_{h}$. We define for any $u_{h},w_{h},v_{h}\in W_{h}^{\mathrm{DG}}$ the following bilinear and trilinear forms:
@@ 4 Analysis of the semi-discrete PolyDG formulation
$\|c\|^{2}_{\mathrm{DG}}=$ $\left\|\sqrt{\mathbf{D}}\nabla_{h}c\right\|^{2}+\left\|\sqrt{\gamma_{F}}[\![c]\!]\right\|^{2}_{\mathscr{F}^{\mathrm{I}}_{h}\cup\mathscr{F}_{h}^{\mathrm{D}}}\qquad\forall c\in W_{1},$ (8a)
$|\!\!\>|\!\!\>|{c}|\!\!\>|\!\!\>|_{\mathrm{DG}}^{2}=$ $\|c\|_{\mathrm{DG}}^{2}+\left\|\gamma_{F}^{-1/2}\{\!\!\{\mathbf{D}\nabla_{h}c\}\!\!\}\right\|_{\mathscr{F}^{\mathrm{I}}_{h}\cup\mathscr{F}_{h}^{\mathrm{D}}}^{2}\qquad\forall c\in W_{2},$ (8b)
where $W_{1}$ and $W_{2}$ are the broken Sobolev spaces defined as Equation (4). We recall that the previous norms are equivalent in the space of discontinuous functions $W_{h}^{\mathrm{DG}}$, under Assumption 2 thanks to the following trace-inverse inequality [29]:
$\|c\|^{2}_{L^{2}(\partial K)}\lesssim\frac{p^{2}}{h_{K}}\|c\|^{2}_{L^{2}(K)}\quad\forall c\in\mathbb{P}_{p}(K)\quad\forall K\in\mathscr{T}_{h}.$ (9)
@@
$\|c(t)\|_{\varepsilon}^{2}=\|c(t)\|^{2}+\int_{0}^{t}\left(\|c(s)\|_{\mathrm{DG}}^{2}+\|c(s)\|^{2}\right)\mathrm{d}s,\qquad|\!\!\>|\!\!\>|c(t)|\!\!\>|\!\!\>|_{\varepsilon}^{2}=\|c(t)\|^{2}+\int_{0}^{t}\left(|\!\!\>|\!\!\>|c(s)|\!\!\>|\!\!\>|_{\mathrm{DG}}^{2}+\|c(s)\|^{2}\right)\mathrm{d}s.$ (10)
@@ 4.1 Preliminary technical results
Let $a,b$ be two positive constants, let $\eta>1$, and let $u\in L^{\infty}_{+}(0,\hat{t})=\{u\in L^{\infty}(0,\hat{t}):\;u(t)\geq 0\;a.e.\,\}$ such that:
$u(t)\leq a+b\int_{0}^{t}u^{\eta}(s)\mathrm{d}s,\qquad\mathrm{for\;almost\;any}\;t\in(0,\hat{t}),$ (11)
@@
$u(t)\leq\dfrac{a}{\left(1-a^{\eta-1}\,b\,(\eta-1)t\right)^{\frac{1}{\eta-1}}}.$ (13)
@@ 7 Simulation of $\alpha$-synuclein spreading in a two dimensional brain section
$\alpha=k_{12}\frac{k_{0}}{k_{1}}-\tilde{k}_{1}.$ (22)